# A Survey of the Self Supervised Learning Mechanisms for Vision Transformers

### v1

@@ 7.1.1 Traditional Augmentation
Augmentation using geometric transformation is being used for various image processing applications. By employing this transformation, the original positions of the image pixels are relocated to new positions without modifying intensity values. It modifies the training data by incorporating variations in viewpoint, non-rigid deformations, and scaling for better model learning on real world problems. The geometric transformations including the translation, rotation, reflection, shearing, and scaling are being commonly used for data augmentation. It is categorized into affine and non-affine transformations. The former one utilized linear mapping for adjusting the geometric disorder in the structure of the image. Whereas the later involved complex non-linear mapping functions [83, 84]. The projective and perspective transformations are fall into the non-affine geometric transformations. These are effectively being used in computer vision and medical imaging tasks [85, 86]. Similarly, the deformation transformations are non-affine which are used to simulate at higher degree of freedom such as Lens and non-rigid body deformations [87, 88, 89]. Photometric transformation is also fall under the traditional augmentation category and used for various computer vision tasks. It considers the camera and shooting artifacts such as optical noise, motion blur, image color artifacts degradations. These simulations are being used to mitigate the effects of data scarcity during model training.
@@ 7.1.2 Advanced Augmentation Strategies
The advanced augmentation strategies are used to solve complex computer vision problems and gained a significant attention of researchers. These augmentation operations are learned from the given data automatically. Several CNN based models are proposed in literature to perform various advanced transformations for data augmentation [90, 91, 92]. These deep models have several cascading layers to learn the transformations, consequently obtained a better trained and generalized model. Jagerberg et al. (2015) proposed spatial transformer network (STN) data augmentation approach which is used for model training. Based on STN data augmentation model, several improved models are built which offer promising results [93, 94, 95]. Although the advanced models are computationally expensive compared to the traditional methods. However, the accuracy and diversity of data augmentation of these models are very high. They are effective to solve various data scarcity related computer vision and SSL problems.

### v2

@@ 7.1.1 Traditional Augmentation
Augmentation using geometric transformation is being used for various image processing applications. By employing this transformation, the original positions of the image pixels are relocated to new positions without modifying intensity values. It modifies the training data by incorporating variations in viewpoint, non-rigid deformations, and scaling for better model learning on real world problems. The geometric transformations including the translation, rotation, reflection, shearing, and scaling are being commonly used for data augmentation. It is categorized into affine and non-affine transformations. The former one utilized linear mapping for adjusting the geometric disorder in the structure of the image. Whereas the later involved complex non-linear mapping functions [109, 110]. The projective and perspective transformations are fall into the non-affine geometric transformations. These are effectively being used in computer vision and medical imaging tasks [111, 112]. Similarly, the deformation transformations are non-affine which are used to simulate at higher degree of freedom such as Lens and non-rigid body deformations [113, 114, 115]. Photometric transformation [116, 117] is also fall under the traditional augmentation category and used for various computer vision tasks. It considers the camera and shooting artifacts such as optical noise, motion blur, image color artifacts degradations. These simulations are being used to mitigate the effects of data scarcity during model training.
@@ 7.1.2 Advanced Augmentation Strategies
The advanced augmentation strategies are used to solve complex computer vision problems and gained a significant attention of researchers. These augmentation operations are learned from the given data automatically. Several CNN based models are proposed in literature to perform various advanced transformations for data augmentation [118, 119, 120]. These deep models have several cascading layers to learn the transformations, consequently obtained a better trained and generalized model. Jagerberg et al. (2015) [118] proposed spatial transformer network (STN) data augmentation approach which is used for model training. Based on STN data augmentation model, several improved models are built which offer promising results [121, 122, 123]. Although the advanced models are computationally expensive compared to the traditional methods. However, the accuracy and diversity of data augmentation of these models are very high. They are effective to solve various data scarcity related computer vision and SSL problems.